If the effective rate of interest is compounded quarterly is 16%, then the nominal rate of interest is


1
Expert's answer
2022-02-10T04:17:30-0500

Answer

Where i = I/100 and r = R/100; nominal interest rate per period

r = m × [ ( 1 + i)1/m - 1 ]

i is the effective rate,

m is the number of compounding periods.

i=0.16

m=4

"r=4\u00d7[(1+0.16)^\u00bc-1]"

r=0.1512

Nominal rate of interest R=15.12%